40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 22 11:33:20 2016, gautam, Update, Green Locking, AUX-X AM/PM investigations AUX_X_AM.pdfAUX_X_RIN.pdfNE_Mephisto.png
    Reply  Fri Jan 22 13:33:37 2016, ericq, Update, Green Locking, AUX-X AM/PM investigations 
       Reply  Fri Jan 22 17:22:06 2016, gautam, Update, Green Locking, AUX-X AM/PM investigations AM_response.pdfNE_investigations.pdf
          Reply  Fri Jan 22 18:46:03 2016, rana, Update, Green Locking, AUX-X AM/PM investigations 
Message ID: 11944     Entry time: Fri Jan 22 11:33:20 2016     Reply to this: 11945
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: AUX-X AM/PM investigations 

I was trying to characterize the AM/PM response of the X end laser. I tried to measure the AM response first, as follows:

  • I used the Thorlabs PDA 55, whose datasheet says it has 10MHz bandwidth - I chose it because it has a larger active area than the PDA 255, but has sufficient bandwidth for this measurement. 
  • My earlier measurement suggested the IR power coming out of the laser is ~300mW. As per the datasheet of the PDA 55, I expect its output to be (1.5 x 10^4 V/A) * (~0.25 A/W) ~ 4000 V/W => I expect the PD output (driving the 50ohm input of the Agilent NA) to saturate at ~1.3mW. So I decided to use a (non-absorptive) ND 3.0 filter in front of the PD (i.e. incident power on the PD ~0.3 mW).
  • I measured the AM response (inputA/inputR) by using the RF output from the Agilent analyzer (divided using a mini-circuits splitter half to input R and half to the laser PZT), and the PD output to input A. I set the power of the RF output on the analyzer to 0 dBm. 
  • Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 
  • I also took a measurement of the RIN with no drive to the laser PZT (terminated with a 50ohm terminator) - see Attachment #2. Qualitatively, this looks like the "free-running" RIN curve on the Innolight datasheet (see Attachment #3, the peak seems slightly shifted to the left though), even though the Noise Eater switch on the laser controller front panel is set to "ON". I neglected taking a spectrum with it OFF, I will update this elog once I do (actually I guess I have to take both spectra again as the laser diode and crystal temperatures have since been changed - this data was taken at T_diode = 28.5deg, I_diode = 1.90A, and T_crystal = 47.5 deg). But does this point to something being broken?
  • I was unable to lock the PLL yesterday to measure the PM response, I will try again today.
Attachment 1: AUX_X_AM.pdf  59 kB  | Show | Hide all | Show all
Attachment 2: AUX_X_RIN.pdf  44 kB  | Hide | Hide all | Show all
AUX_X_RIN.pdf
Attachment 3: NE_Mephisto.png  75 kB  | Hide | Hide all | Show all
NE_Mephisto.png
ELOG V3.1.3-